# SVG360: Multi-View SVG Generation with Geometric and Color Consistency from a Single SVG

### v1

@@ Abstract
Scalable Vector Graphics (SVGs) are central to modern design workflows, offering scaling without distortion and precise editability. However, for single object SVGs, generating multi-view consistent SVGs from a single-view input remains underexplored. We present a three stage framework that produces multi-view SVGs with geometric and color consistency from a single SVG input. First, the rasterized input is lifted to a 3D representation and rendered under target camera poses, producing multi-view images of the object. Next, we extend the temporal memory mechanism of Segment Anything 2 (SAM2) to the spatial domain, constructing a spatial memory bank that establishes part level correspondences across neighboring views, yielding cleaner and more consistent vector paths and color assignments without retraining. Finally, during the raster to vector conversion, we perform path consolidation and structural optimization to reduce redundancy while preserving boundaries and semantics. The resulting SVGs exhibit strong geometric and color consistency across views, significantly reduce redundant paths, and retain fine structural details. This work bridges generative modeling and structured vector representation, providing a scalable route to single input, object level multi-view SVG generation and supporting applications such as asset creation and semantic vector editing.
@@ 1 Introduction
Editable vector graphics are central to modern design workflows, providing resolution independence, structural clarity, and precise editability. However, generating multi view consistent and fully editable vector graphics for a given object remains an open challenge. When designers attempt to present an existing vector illustration from a new viewpoint, they often need to manually redraw or adjust its structure to approximate the desired perspective. This manual process is both time consuming and prone to geometric and stylistic inconsistencies. Adobe Illustrator’s Turntable (Beta) [1] is one of the few publicly available end-to-end tools that generate multiple views from a single 2D vector artwork, yet it provides constrained angle control and discloses no technical specification.
@@
The task of automatically generating multi-view consistent and fully editable SVGs, however, has not been extensively explored. We define this task as synthesizing a set of path level editable SVGs that preserve geometric and stylistic coherence across viewpoints. Solving this problem is particularly valuable for creative workflows such as turntable style visualization, multi-view logo and icon generation, and geometry consistent SVG asset editing. Bridging the gap between generative modeling and structured vector representation is therefore critical for enabling scalable and reliable design pipelines.
@@
Recent advances in generative AI, particularly diffusion models[10, 29, 25, 30, 26, 46, 23], have enabled photorealistic and view-consistent image synthesis from a single input view, with representative approaches including a series of image-conditioned multi-view generation models [20, 33, 21, 38].
@@
Yet, these outputs remain raster based and lack the structured semantics and fine grained editability inherent to vector graphics. Meanwhile, existing vectorization approaches such as DiffVG [18], DeepSVG [3], and Im2Vec [28] focus on single view reconstruction and fail to maintain geometric consistency across views. This gap motivates a core objective: enabling generative models to produce multi view coherent and structurally editable representations directly in the SVG domain.
@@
To realize this objective, we propose a unified generative to vector framework that produces multi-view, geometrically consistent, fully editable SVGs from a single input. Our method consists of three stages. It first performs three dimensional aware multi-view generation to obtain geometry plausible novel views. Then it performs cross-view refinement through spatially aligned segmentation and appearance harmonization, enhancing the consistency of object contours and colors across views in the SVG style domain. Finally, leveraging the multi-view consistent segmentation results, we perform raster-to-vector conversion followed by path consolidation and color correction, yielding compact, coherent, and easily editable multi-view SVGs.
On multi-view SVG generation, our method produces noticeably cleaner and more stable results than existing vectorization tools. Compared with Adobe Illustrator’s Turntable, the path counts of our generated views stay much closer to the structure of the input SVG, reducing the path-count deviation from the input SVG by 26.5%. The variation in color usage between adjacent viewpoints is also reduced by 83.5%, indicating a far more consistent preservation of the input color design across rotations. In addition, our method lowers the average number of paths by 11.6%, resulting in lighter and more editable multi-view SVG sequences. The generated views exhibit coherent appearance transitions and show strong potential for creative design workflows, turntable-style visualization, and semantically aligned vector editing. Our contributions are summarized as follows:
@@
Unified Generative to Vector Framework. We introduce the first end-to-end framework to directly synthesize multi view consistent and fully editable SVGs from a single input, bridging generative modeling with structured vector representation.
@@
Spatially-Aware Cross-View Segmentation Refinement. We design a spatially-aware segmentation propagation mechanism that replaces temporal adjacency with geometry-guided neighborhood traversal, ensuring consistent part-level segmentation and enhancing SVG generation accuracy across multiple views.
@@ 2 Related Work
2.1 Single image to multi view generation
@@
Turning one reference view into a usable turntable sequence has become a key precursor to our goal of editable multi-view SVGs. Existing approaches fall into two major categories:
@@
Diffusion-based multi-view image synthesis builds on high-fidelity 2D diffusion priors. These methods condition the denoising trajectory on target camera poses to synthesize each view directly in the image domain. Representative works include Zero-1-to-3 [20] and Zero123++ [33], where the former targets novel-view images rather than explicit 3D, and the latter improves cross-view consistency through better conditioning and training. Further, SyncDreamer [21] explicitly couples views during sampling via synchronized diffusion to align geometry and color, and SV3D [38] leverages video-latent diffusion to generate orbital view sequences from a single image. Despite strong appearance fidelity, residual independence across views often leads to geometric and color drift on thin structures and self occlusions.
@@
3D reconstruction and generation from single images. These methods predict a 3D asset from one or a few views and render all novel viewpoints from the shared geometry, which naturally improves multi-view geometric consistency over 2D diffusion synthesis. Representative approaches include mesh generators such as GET3D, PolyGen, MeshDiffusion [8, 24, 22], single-image reconstruction methods like LRM, TripoSR, UniQue3D [11, 36, 40], and latent 3D models including Shape-E and Trellis [14, 41]. However, their outputs remain rasterized images without explicit vector paths or cross-view path correspondences, making them unsuitable for SVG-oriented workflows.
3D Gaussian Splatting (3DGS) as a shared representation. 3DGS [15] represents a scene with anisotropic Gaussians and renders by fast differentiable splatting, so all views reproject the same underlying 3D field. This often yields strong cross view geometric stability and efficient multi view rendering. Recent works combine diffusion priors with 3DGS for rapid 3D asset creation [35, 7], and single image variants enable ultra fast inference [34]. Although 3DGS provides efficient and geometrically consistent multi-view rendering, its outputs remain raster-based and do not explicitly expose vector path topology. Consequently, converting rendered contours into clean, closed, and editable paths requires an additional vectorization stage.
2.2 Raster to Vector Generation
Transforming multi-view consistent rasters into structured vector representations remains challenging because recovering continuous topology and geometry from discrete pixels is inherently ill-posed. Classical contour tracing methods such as Potrace and AutoTrace detect boundaries and fit Bézier curves or polygons. They perform well on icons or low-color imagery, but on smooth gradients and complex textures they often produce jagged, fragmented, or redundant paths.
Learning-based approaches model vector paths parametrically. DiffVG [18] introduces a differentiable renderer for gradient-based optimization, and DeepSVG [3] and Im2Vec [28] leverage Transformer or VAE designs to improve geometric continuity and structural controllability. LIVE [45] decomposes images into hierarchical layers to preserve global topology, yet can over-segment fine details. More recent methods such as VectorFusion [13] and StarVector [39] incorporate diffusion priors or region fusion to obtain smoother and more coherent paths.
@@ 2.2 Raster to Vector Generation
Despite these advances, all vectorization methods still operate directly on pixel-level color transitions, which makes complex objects prone to fragmented or redundant paths. Introducing semantic segmentation before vectorization greatly simplifies the problem by isolating coherent part regions, enabling cleaner per-part conversion and more stable geometric structures. This highlights the need for a semantic layer, especially when extending vectorization to multi-view settings.
2.3 Segmentation and Cross View Consistency
@@
While semantic segmentation simplifies single-view vectorization by isolating coherent part regions, multi-view SVG generation raises an additional requirement: achieving more consistent part assignments across viewpoints greatly improves cross-view coherence and visual stability. Yet conventional segmentation relies mainly on per-image pixel cues, making such consistency difficult; a spatially stable segmentation step is therefore needed before vectorization to provide more reliable part structures across views.
@@
Early two-dimensional methods [9, 4, 42, 17] focus on per-image quality but do not enforce label correspondence across different views of the same object, causing boundaries to drift or merge inconsistently after projection. Foundation-scale models such as SAM, HQ-SAM, and SAM2 [16, 5, 27] provide accurate boundaries and promptable interaction, thus serving as strong building blocks for multi-view scenarios. However, their propagation mechanisms are primarily temporal rather than spatial, which limits their ability to maintain consistent part correspondences across viewpoints in static 3D scenes.
@@
Inspired by video object segmentation, memory-based propagation [6, 43] maintains features across frames to keep masks consistent over time. By analogy, multi-view consistency can be formulated as a spatial sequence problem on the viewing sphere, where each camera view is treated as a pseudo-temporal step and neighboring views act as spatial references.
@@
Existing work either stabilizes masks in raster space or improves single-view vectorization quality, but a bridge between cross-view semantic consistency and vector-level path correspondence remains underexplored. This gap motivates our approach, which first achieves part-aware, cross-view consistent segmentation and then leverages these semantic units to drive compact and editable vector representations across views.
@@ 3 Method
Figure 1: Our pipeline begins by converting the input SVG into a raster image, followed by rendering 3D consistent multi-view rasters using Trellis. A lightweight LoRA [12] tuned FLUX [2] model is then applied to harmonize their appearance. The refined rasters are processed by our Spatial SAM2 module, which replaces temporal adjacency in SAM2 with a spatial nearest-neighbor traversal on the viewing sphere. During segmentation propagation, the Spatial Memory Selector retrieves the geometrically most relevant memory entries. For the target view $t_{n-3}$, spatially adjacent views such as $t_{9}$, $t_{55}$, or $t_{n-6}$ may be closer than its temporal neighbor $t_{n-4}$, which helps maintain part-level consistency across viewpoints. The resulting masks are vectorized by VTracer [37] and further refined in the vector domain to produce compact, editable, and cross-view consistent multi-view SVGs.
@@ 3.1 Multi-View Raster Generation and Harmonization
Multi-view raster generation. Generating accurate and geometrically consistent multi-view images from a single SVG input is challenging, as purely 2D generation approaches struggle to ensure structural coherence across views. To address this, we adopt a generative 3D-based approach to synthesize multi-view images.
Although several recent single-image 3D methods (e.g., Sparc3D [19] and Hi3DGen [44]) demonstrate strong geometric fidelity, they primarily focus on precise shape reconstruction and often do not provide fully textured, render ready outputs. Such geometry-only representations are unsuitable for our color consistent raster to SVG multi-view workflow. Considering open-source availability, generation efficiency, and the level of visual fidelity required by our task, we adopt Trellis [41] as the backbone 3D generator. Trellis offers a balanced trade-off between geometric accuracy and texture consistency, and remains efficient and practical for producing coherent multi-view rasters from a single SVG input.
Trellis represents assets in a unified structured latent space and decodes them into multiple 3D formats, including meshes and 3D Gaussians [41]. Its mesh decoder builds on FlexiCubes [32], predicting per-voxel signed-distance values and extracting a watertight surface from the zero-level isosurface. While both decoders yield high-fidelity results, converting the latent to a mesh and then rasterizing from many viewpoints introduces extra discretization and baking steps, which may accumulate appearance deviations relative to the input SVG style and incur nontrivial runtime overhead. For efficiency and fidelity, we directly use the 3D Gaussian decoder and render dense multi-view rasters via Gaussian splatting, which preserves the appearance statistics needed by our downstream segmentation and vectorization stages.
@@
While a 3D proxy preserves global geometric consistency across viewpoints, the reconstruction can still exhibit local errors, especially along edges and curved surfaces. View dependent illumination further introduces color variation that destabilizes raster-to-SVG conversion. We apply appearance harmonization using FLUX.1-dev [2] fine-tuned with a LoRA [12] trained on flat-design SVG data. The harmonized renders increase intra-region homogeneity and boundary contrast, suppress spurious shading edges, and reduce color variation across views. The resulting rasters are both view-consistent and faithful to SVG-style appearance, which improves downstream segmentation separability and stabilizes subsequent vector path consolidation.

### v2

@@ Abstract
Scalable Vector Graphics (SVGs) are central to modern design workflows, offering scaling without distortion and precise editability. However, for single object SVGs, generating multi-view consistent SVGs from a single-view input remains underexplored. We present a three stage framework that produces multi-view SVGs with geometric and color consistency from a single SVG input. First, the rasterized input is lifted to a 3D representation and rendered under target camera poses, producing multi-view images of the object. Next, we extend the temporal memory mechanism of Segment Anything 2 (SAM2) to the spatial domain, constructing a spatial memory bank that establishes part-level correspondences across neighboring views, yielding cleaner and more consistent vector paths and color assignments without retraining. Finally, during the raster-to-vector conversion, we perform path consolidation and structural optimization to reduce redundancy while preserving boundaries and semantics. The resulting SVGs exhibit strong geometric and color consistency across views, significantly reduce redundant paths, and retain fine structural details. This work bridges generative modeling and structured vector representation, providing a scalable route to single input, object level multi-view SVG generation and supporting applications such as asset creation and semantic vector editing.
@@ 1 Introduction
Editable vector graphics are central to modern design workflows, providing resolution independence, structural clarity, and precise editability. However, generating multi-view consistent and fully editable vector graphics for a given object remains an open challenge. When designers attempt to present an existing vector illustration from a new viewpoint, they often need to manually redraw or adjust its structure to approximate the desired perspective. This manual process is both time consuming and prone to geometric and stylistic inconsistencies. Adobe Illustrator’s Turntable (Beta) [adobeTurntableHelp] is one of the few publicly available end-to-end tools that can generate multiple views from a single 2D vector artwork. However, it offers limited control over viewing angles and discloses no technical specifications.
@@
The task of automatically generating multi-view consistent and fully editable SVGs, however, has not been extensively explored. We define this task as synthesizing a set of path-level editable SVGs that preserve geometric and stylistic coherence across viewpoints. Solving this problem is particularly valuable for creative workflows such as turntable style visualization, multi-view logo and icon generation, and geometry consistent SVG asset editing. Bridging the gap between generative modeling and structured vector representation is therefore critical for enabling scalable and reliable design pipelines.
@@
Recent advances in generative AI, particularly diffusion models [ho2020ddpm, rombach2022ldm, nichol2021glide, saharia2022imagen, ramesh2022dalle2, zhang2023controlnet, mou2023t2iadapter], have enabled photorealistic and view-consistent image synthesis from a single input view, with representative approaches including a series of image-conditioned multi-view generation models [liu2023zero123, shi2023zero123pp, liu2023syncdreamer, voleti2024sv3d].
@@
Yet, these outputs remain raster-based and lack the structured semantics and fine grained editability inherent to vector graphics. Meanwhile, existing vectorization approaches such as DiffVG [li2020diffvg], DeepSVG [carlier2020deepsvg], and Im2Vec [reddy2021im2vec] focus on single-view reconstruction and fail to maintain geometric consistency across views. This gap motivates a core objective: enabling generative models to produce multi-view coherent and structurally editable representations in the SVG domain.
@@
To realize this objective, we propose a unified generative to vector framework that produces multi-view, geometrically consistent, fully editable SVGs from a single input. Our method consists of three stages. It first performs 3D-aware multi-view generation to obtain geometrically plausible novel views. Then it performs cross-view refinement through spatially aligned segmentation and appearance harmonization, enhancing the consistency of object contours and colors across views in the SVG style domain. Finally, leveraging the multi-view consistent segmentation results, we perform raster-to-vector conversion followed by path consolidation and color correction, yielding compact, coherent, and easily editable multi-view SVGs.
On multi-view SVG generation, our method produces noticeably cleaner and more stable results than a strong industrial baseline. Compared with Adobe Illustrator’s Turntable, our path count deviation ($\mathrm{RMSE}_{\text{path}}$) is about 23% lower, and the adjacent-view color count variation ($\overline{\Delta N}_{\text{color,nbr}}$) is about 86% lower. We also generate more compact SVGs, with about 7% fewer paths and about 17% fewer colors on average. In addition, our method yields smoother transitions on the yaw-pitch grid, with around 3% higher DINO feature consistency and around 2.5% higher DreamSim perceptual similarity on pitch changes [oquab2023dinov2, fu2023dreamsim]. Finally, our vectorization module achieves high single-view fidelity measured by SSIM, LPIPS, and $\Delta E_{2000}$ [wang2004image, zhang2018unreasonable, sharma2005ciede2000]. On our single-view benchmark, it attains the lowest $\Delta E_{2000}$ among the compared raster-to-SVG baselines while maintaining a compact topology [yue2024live, wang2024starvector, weber2001autotrace, jain2022vectorfusion].
@@
Unified Generative to Vector Framework. We introduce an end-to-end framework to directly synthesize multi-view consistent and fully editable SVGs from a single input, bridging generative modeling with structured vector representation.
@@
Spatially Aware Cross-View Segmentation Refinement. We design a spatially aware segmentation propagation mechanism that replaces temporal adjacency with geometry-guided neighborhood traversal, ensuring consistent part-level segmentation and enhancing SVG generation accuracy across multiple views.
@@ 2 Related Work
2.1 Single-Image to 3D Vector Sketches
Such methods predict sparse 3D contour curves from single-view renders [choi2024_3doodle, li2025_empowering, wang2025_viewcraft3d]. Lacking filled regions and explicit colors, they inevitably lose the original SVG style. Instead, we desire to generate multi-view 2D SVGs that strictly preserve region geometry and design semantics.
2.2 Single-image to multi-view generation
@@
Generating a turntable-like view sequence from a single reference image is a common precursor to editable multi-view SVGs. Existing approaches can be grouped into three categories.
@@
Diffusion-based multi-view image synthesis. These methods condition 2D diffusion models on target camera poses to synthesize views in the image domain. Representative works include Zero-1-to-3 and Zero123++ [liu2023zero123, shi2023zero123pp], SyncDreamer [liu2023syncdreamer], and SV3D [voleti2024sv3d]. Despite strong appearance fidelity, view-wise sampling often causes geometric and color drift on thin structures and self-occlusions.
@@
Single-image 3D reconstruction and generation. These methods infer a shared 3D asset and render novel views from the same geometry, improving multi-view consistency over 2D diffusion. Representative approaches include mesh generators (GET3D, PolyGen, MeshDiffusion) [gao2022get3d, polygen, meshdiffusion], single-image reconstruction (LRM, TripoSR, UniQue3D) [hong2023lrm, triposr, unique3d], and latent 3D models (Shape-E, Trellis) [shapee, trellis2024]. However, the outputs are still raster images without explicit vector paths or cross-view path correspondences.
3D Gaussian Splatting (3DGS) as a shared representation. 3DGS [kerbl2023gaussians] enables efficient rendering by reprojecting a shared 3D field, often yielding stable multi-view geometry. Recent works combine diffusion priors with 3DGS for rapid asset creation [tang2023dreamgaussian, gaussiandreamer2023] and extend to single-image variants [splatterimage2024]. Nevertheless, the results remain raster-based and require an additional vectorization stage to obtain clean, closed, and editable SVG paths.
2.3 Raster-to-Vector Generation
Converting rendered rasters into structured, editable vectors is challenging, especially when consistency must be maintained across multiple views. Classical tracing methods such as Potrace [selinger2003potrace] and AutoTrace [weber2001autotrace] work well for icons and low-color graphics, but often produce fragmented or redundant paths under smooth shading and complex textures.
Prior work includes optimization-based methods, e.g., DiffVG [li2020diffvg] which optimizes control points via differentiable rendering, and learning-based methods such as DeepSVG [carlier2020deepsvg] and Im2Vec [reddy2021im2vec] built on Transformer [vaswani2017attention] or VAE [kingma2014vae] designs. For raster-to-SVG conversion, LIVE [yue2024live] preserves global topology via hierarchical layers but can over-segment fine structures, while VectorFusion [jain2022vectorfusion] and StarVector [wang2024starvector] incorporate diffusion priors or region-aware fusion for smoother paths.
@@ 2.3 Raster-to-Vector Generation
Despite these advances, many pipelines still rely heavily on local pixel-level color transitions, making complex objects prone to over-fragmentation in multi-view settings. Introducing part-level segmentation before vectorization isolates coherent regions, enabling cleaner per-part conversion and more stable structure and palette behavior across views.
Figure 1: Our pipeline begins by converting the input SVG into a raster image, followed by rendering 3D consistent multi-view rasters using Trellis. We then apply FLUX.1 dev [blackforestlabs2024flux] with a public vector-style LoRA [hu2021lora] adapter (SimpleVectorFlux [renderartist2024simplevectorflux]) to harmonize their appearance. The refined rasters are processed by our Spatial SAM2 module, which replaces temporal adjacency in SAM2 with a spatial nearest-neighbor traversal on the viewing sphere. During segmentation propagation, the Spatial Memory Selector retrieves the viewpoint-wise nearest memory entries. For the target view $t_{n-3}$, spatially adjacent views such as $t_{9}$, $t_{55}$, or $t_{n-6}$ may be closer than its temporal neighbor $t_{n-4}$, which helps maintain part-level consistency across viewpoints. The resulting masks are vectorized by VTracer [vtracer2023] and further refined in the vector domain to produce compact, editable, and cross-view consistent multi-view SVGs.
2.4 Segmentation and Cross-view Consistency
@@
While semantic segmentation simplifies single-view vectorization by isolating coherent part regions, multi-view SVG generation additionally requires consistent part assignments across viewpoints. However, conventional segmentation relies on per-image pixel cues, making such consistency difficult; we therefore need a spatially stable segmentation step before vectorization to provide reliable part structures across views.
@@
Early two-dimensional methods [he2017maskrcnn, cheng2022mask2former, xie2021segformer, li2023maskdino] focus on per-image quality but do not enforce label correspondence across different views of the same object, causing boundaries to drift or merge inconsistently after projection. Foundation-scale models such as SAM, HQ-SAM, and SAM2 [kirillov2023sam, cheng2023hqsam, ravi2024sam2] provide accurate boundaries and promptable interaction, but their propagation mechanisms are primarily temporal rather than spatial, which limits consistent part correspondences across viewpoints in static 3D scenes.
@@
Memory-based propagation in video object segmentation [cheng2021stcn, yang2023deAOT] maintains features across frames to keep masks consistent over time. In multi-view settings, we can similarly view cameras as a spatial sequence on the viewing sphere, treating each view as a pseudo-temporal step and using neighboring views as spatial references.
@@
A bridge between cross-view semantic consistency and vector-level path correspondence remains underexplored. Our approach addresses this gap by first achieving part-aware, cross-view consistent segmentation, and then leveraging these semantic units to drive compact and editable vector representations across views.
@@ 3.1 Multi-View Raster Generation and Harmonization
Multi-view raster generation.
Generating accurate and geometrically consistent multi-view images from a single SVG input is challenging, as purely 2D generation approaches struggle to ensure structural coherence across views. To address this, we adopt a generative 3D-based approach to synthesize multi-view images.
Although several recent single-image 3D methods (e.g., Sparc3D [sparc3d2025] and Hi3DGen [hi3dgen2025]) demonstrate strong geometric fidelity, they primarily focus on precise shape reconstruction and often do not provide fully textured, render ready outputs. Such geometry-only representations are unsuitable for our color consistent raster to SVG multi-view workflow. Considering open-source availability, generation efficiency, and the level of visual fidelity required by our task, we adopt Trellis [trellis2024] as the backbone 3D generator. Trellis offers a balanced trade-off between geometric accuracy and texture consistency, and remains efficient and practical for producing coherent multi-view rasters from a single SVG input.
Trellis represents assets in a unified structured latent space and decodes them into multiple 3D formats, including meshes and 3D Gaussians [trellis2024]. Its mesh decoder builds on FlexiCubes [flexicubes2023], predicting per-voxel signed-distance values and extracting a watertight surface from the zero-level isosurface. While both decoders yield high-fidelity results, converting the latent to a mesh and then rasterizing from many viewpoints introduces extra discretization and baking steps, which may accumulate appearance deviations relative to the input SVG style and incur nontrivial runtime overhead. For efficiency and fidelity, we directly use the 3D Gaussian decoder and render dense multi-view rasters via Gaussian splatting, which preserves the appearance statistics needed by our downstream segmentation and vectorization stages.
@@
While a 3D proxy preserves global geometric consistency across viewpoints, the reconstruction can still exhibit local errors, especially along edges and curved surfaces. View-dependent illumination further introduces color variation that destabilizes raster-to-SVG conversion. We apply appearance harmonization using FLUX.1-dev [blackforestlabs2024flux] with a public vector-style LoRA adapter (SimpleVectorFlux [renderartist2024simplevectorflux]). The harmonized renders increase intra-region homogeneity and boundary contrast, suppress spurious shading edges, and reduce color variation across views. The resulting rasters are both view-consistent and faithful to a flat vector appearance, which improves downstream segmentation separability and stabilizes subsequent vector path consolidation.